OpenStudy (anonymous):

Write a linear system that models this situation. 1.You have a total of 5 coins.Some are nickels some are dime. the total value is 40 cents. 2. 5 bagels and 4muffins cost $7 together. a muffin cost 0.40 more than the bagel.

OpenStudy (anonymous):

x+y=5 0.05x+0.10y=0.40

OpenStudy (anonymous):

can you help me on #2 too~~please

OpenStudy (anonymous):

5x+4y=7 y=0.40+x